OpenStudy (anonymous):

A square has an area of 256 cm^2. What is the length of one side of the square? Answer (A) 128 cm (C) 64 cm (D) 16 cm (E) 4 cm

OpenStudy (anonymous):

because square have 4 sides

OpenStudy (goformit100):

(C) 64 cm

OpenStudy (campbell_st):

the area is \[A = (length)^2 ... or... A = l^2\] so the opposite is taking the square root so \[256 = l^2 .... then.... \sqrt{256} = l\]

OpenStudy (anonymous):

Thanks You guys :*